Project

General

Profile

News

Update 7 August 2017

Added by Kyberd, Paul 2 months ago

Update: liquid-hydrogen system commissioning:

The third and final test of the liquid-hydrogen system using neon was successful. The test ended last Thursday (03Aug17) and the system was allowed to warm up over the weekend. The warm-up has been slow, indicating excellent thermal installation.

The outstanding work that must be carried out before the final safety inspection includes the completion of the remedial work on the hydrogen-quench line, the installation and inspection of ATEX-rated fittings in the LH2 shed on the roof of then MICE Hall and the completion of the safety-related paper work. In addition, the leak testing in the hydrogen-gas panel has to be repeated to document the (low and satisfactory) leak rates. The liquid-hydrogen team is working steadily to address the last remaining issues.

The final safety tour is scheduled for 17Aug17. Assuming the system is signed off as safe to operate, the commissioning of the system with hydrogen will commence as soon after the tour as possible.

New public results for upcoming conferences, e.g., COOL17 and NuFact17:

Progress on our various analyses is good and the analysis teams are pushing hard to complete the work necessary to bring forward new results. Two conferences that are important for us to present our work are on the horizon, COOL17 (18—22 September 2017) and NuFact17 (25–30 September 2017).

In view of the approaching conferences, I thought it timely to restate the procedure by which we make new results public. The results must be presented to the collaboration at a video conference or collaboration meeting so that the results can be discussed and agreed before they are made public. An entry is then made on the public information page to document the new results.

For the upcoming conferences, the video conference on the 7th September 2017 will be the time when new results should be presented to the collaboration. The 7th is perfectly timed for COOL17 as it is two weeks ahead of the conference. Since COOL17 and NuFact17 are only one week apart, the video conference on the 7th will serve for NuFact as well.

With best wishes …

Ken

Update 25 July 2017

Added by Kyberd, Paul 2 months ago

Update: liquid-hydrogen system commissioning:

Progress on the preparation of the liquid-hydrogen system has been steady since CM48 in Belgrade. The modifications to the turret required to address the issues raised during commissioning with neon have all been addressed. The turret has been installed in the focus coil and the connections to the liquid-hydrogen vessel have been made.

Last week, the absorber/focus-coil module was placed back on the beam line. Vacuum in the hydrogen vessel, the safety volume and the interspace was established at the required levels and the cool-down of the system with helium gas was initiated by Mark Tucker on Saturday 22Jul17, a milestone for the preparation of the system. Another milestone was passed yesterday (Monday, 24Jul17) morning when the temperature of the vessel fell below 75K; the temperature at which the change is made from helium to the gas which will be liquified, in this case neon. Neon is presently being liquified by the system; the third, and final, test of the system with neon is now underway.

On Thursday 20Jul17, the preliminary safety tour of the liquid-hydrogen system took place. This is the first formal step in gaining permission to operate the system with liquid hydrogen. It is testament to the care with which the liquid-hydrogen team have prepared the system that only a few, minor, issues were identified by the reviewers. These issues will be addressed once the final system test with neon is complete.

The timetable now is:
• A full system test with liquid neon will be performed. This test will include the test of each of the sequences required during operation with hydrogen. The safety systems will also be tested. It is anticipated that the full set of tests will complete around 04Aug17;
• The system will be warmed up to allow the hydrogen-quench line to be finalised. In parallel to the completion of the hydrogen-quench line, a final round of leak checking on the hydrogen-gas panel will be carried out. It is anticipated that this will take place in the week of 07Aug17;
• Following the completion of the above work, a second, hopefully final, safety tour will take place that, if successful, will allow the system to be signed off as safe to operate with hydrogen.

The culmination of the steps listed above, the second hydrogen safety tour, should take place in around three weeks time. I would like to congratulate Josef Boehm, Mark Tucker, Phil Warburton and the whole of the liquid-hydrogen team for pushing the work past the two critical milestones that have just been achieved.

Preparation for operation in Cycle 2017/02:

The next User Cycle, 2017/02 starts on 19Sep17, in only eight-weeks time. Steve Boyd and Colin Whyte will take the lead in the coordination of the steps necessary to bring up all the MICE systems. Careful scheduling will be required to dovetail the work on the superconducting magnets, the tracker readout, the MICE Muon Beam and target and all the instrumentation, readout and control systems that we require to run.

Of particular importance in the short term is a thorough and careful review of the alarm handler, its limits and the way in which it is used. This will be coordinated by the Duty Coordinator (presently Paul Hodgson). Ajit Kurup will implement necessary changes to alarm limits etc. The “owners” of each system giving an alarm will be contacted and supported to make the changes necessary to remove the alarm or to perform the analysis required to set more appropriate limits. Everyone appreciates the importance of this work, it is a necessary first step to bringing up the system for data taking.

In Belgrade we discussed the programme that we wish to carry out in the coming Cycle and identified the bones of a strong case for extended running. However, we need to be clear that Cycle 2017/02 may be our last. This means that we need to be ready to start when ISIS delivers stable operation on 19Sep17, which in turn means that our pre-Cycle preparations must be both efficient and effective. Plans have been laid for the recommissioning of the magnets, the tracker electronics and readout and other systems. I am confident that we shall be in good shape for data taking in September.

With best wishes …

Ken

Update 26 May 2017

Added by Kyberd, Paul 2 months ago

Update; operations to end of Cycle 2017/01:

ISIS User Cycle 2017/01, presently underway, will end at 08:00 GMT on Friday 02Jun17. The following programme has been established for operations until the end of the Cycle:
• Long-term “soak” test of the LH2 system with liquid neon:
⁃ On Friday 26May17, a small quantity of neon was liquified. The system was set to maintain the neon in the liquid state over the long weekend. If this is successful it will prove the temperature and pressure stabilisation system over a period of 64—72 hours. The system is operating well, with the various temperatures and pressures being maintained within the pre-determined limits.
⁃ Next week further commissioning of the level sensors and the empty, fill and emergency sequences will be performed.
• Data taking:
⁃ Ed Overton (MOM) with Craig Macwaters, Alan Bross and others have re-established stable operation of the tracker cryostats. Yesterday (Saturday) an expert-led start-up, including the tracker readout, was performed. All four tracker cryostats are now at operating temperature and 24/7 data taking started today.
⁃ Data will be taken without magnetic field in the channel for:
⁃ Tracker calibration and alignment;
⁃ Detailed investigation of aperture limitations introduced by the LH2-absorber vessel. This will give us an important head-start on the LH2 scattering and emittance-evolution measurements; and if possible,
⁃ A scattering measurement on cold gaseous neon.

CM48, Belgrade, 27–29 June 2017

It is now only one month to go to the collaboration meeting in Belgrade. Next week we’ll begin to assemble the agenda for the meeting. We will use the meeting to push forward our analyses and make progress towards the publication of the data that we have already made public. If you’d like to make a presentation please drop me a line.

Registration for CM48 is open. Please register at the meeting page:

https://indico.cern.ch/event/626707/

Early registration is important to allow Dimitrije and his colleagues to make the necessary arrangements and to ensure that your choice of accommodation is available!

With best wishes …

Ken

Update 22 May 2017

Added by Kyberd, Paul 2 months ago

New results presented at IPAC17

The MICE expedition to IPAC included contributions from Chris Rogers, Paulo Franchini, Chris Hunt and Tanaz Mohayai:
• Chris R presented the first results of our study of the evolution of single-particle amplitudes (http://micewww.pp.rl.ac.uk/projects/analysis/wiki/Publications_and_Figures_Step_IV#Study-of-Ionization-Cooling-with-the-MICE-Experiment-Preliminary). The goal now is to bring the analysis to the state where we can publish a “rapid communication” by the end of the summer;
• Paulo presented the Step IV emittance-measurement and scattering results. His poster contained a first pair of plots from the energy loss analysis (S. Wylbur, see http://micewww.pp.rl.ac.uk/projects/analysis/wiki/Publications_and_Figures_Step_IV#Energy-Loss-in-Lithium-Hydride-and-Hydrogen-Absorbers-Preliminary);
• Chris H presented the performance of the cooling demonstration configuration accepted for publication in PRAB and the simplified scenario developed as an upgrade to Step IV (see http://micewww.pp.rl.ac.uk/projects/analysis/wiki/Publications_and_Figures_Step_IV#Layout-of-the-MICE-Demonstration-of-Muon-Ionization-Cooling-Including-Simplified-Design); and
• Tanaz presented the MC study of her “kernel-density estimator” analysis.

Progress on the commissioning of the liquid-hydrogen system in the Hall

The commissioning programme for the liquid-hydrogen delivery system in the MICE Hall has been progressing steadily. The system was cooled down to operating temperature last week. To prove the system, the commissioning programme includes the liquefaction of neon using the full LH2 system. This test is now underway. Neon was introduced into the system on Friday. Measurements of system parameters indicate that the cooling power is more than sufficient but the power that can be supplied by the heaters on the second stage of the cryocooler is too low to allow the system to be regulated. The heater power can be increased by exchanging the voltage source that drives current through the heaters. This will be done tomorrow (Monday 22May17).

The immediate programme for the LH2-system commissioning is the verification of the control sequences. This will be carried out between Monday 22May17 and Wednesday 24May17. On Wednesday, the system will be warmed up so that the absorber can be moved to R9 and to allow the beam-line to be configured for data taking.

On the 16th and 17th May 2017, a full HAZOP of the system was carried out in the light of changes made in response to experience gained in operating the LH2 system in R9 and the operation of the superconducting-magnet channel. The HAZOP recommended a number of actions be completed before the system is operated with hydrogen. The necessary modifications will be implemented and the system reinstalled such that commissioning can be completed in June 2017.

Status of the MICE apparatus

The expert-led startup of the experiment on the 08May went well; with the exception of the tracker, all detectors and the beam line operated without major issues.

The tracker cryostats have been a cause of concern. It was decided to carry out routine maintenance on two of the four cold-heads. When “cryo-2” was cooled down after maintenance, a large number of new dead channels were observed in the readout. This is now believed to be due to the freezing of water. While the maintenance operations were underway, the cryocooler on cryo-4 failed and it was decided to warm up the remaining two cryostats to allow maintenance on their cold-heads to be carried out.

In the light of the number of bad channels generated in the cool-down of cryo-2, a careful study of the content of the gas in the insulating vacuum space was carried out. This work culminated when Alan Bross, Ed Overton and Craig Macwaters discovered substantial amounts of water vapour using a custom liquid-nitrogen cold trap that Alan brought to RAL from FNAL. Using the cold trap, a programme to remove the water vapour has been carried out successfully on cryo-3. The programme to “dry out” the remaining cryostats is now nearing completion and the tracker group anticipates that the tracker will be ready for data taking by the end of the week.

The MICE Operations Managers for this Cycle, Victoria Blackmore and Ed Overton, are doing a good and patient job of responding with their usual professionalism and enthusiasm to the challenges that the hardware has been presenting them with!

Plans for data taking in Cycle 2017/01
The shift schedule for the present User Cycle (2017/01) was arranged to culminate in 24/7 running for the last shift block of the Cycle (next week). Given the outcome of the LH2-system HAZOP, the modifications required to the LH2 system and the fact that the tracker will be ready for operation at the end of the coming week, the run plan has been revisited to optimise the use of the data-taking time.

Analysis of the emittance-evolution data sets indicates that careful studies will be required to understand the transmission through the cooling cell. In particular, we need to study carefully the apertures that limit the transmission or cause energy-loss and scattering that may bias our measurements. After consultation and some discussion, the Analysis Group has requested empty-channel data in the magnet configurations in which the emittance-evolution data was taken. This data will allow us to study transmission in the absence of the effect of the LiH absorber.

In the light of the considerations outlined above, the programme for the rest of the present Cycle will be:
• Continued evaluation of the LH2 delivery system with emphasis on proving the controls sequences: Monday (22May17)—Wednesday (24May17);
• Extraction of LH2 absorber, re-insertion of the FC and establishment of the magnetic channel: Thursday (25May17)—Saturday (27May17);
• Magnet commissioning: Sunday (28May17)—Monday (29May17); and
• Empty-channel data taking (Monday 29May17)—Friday (02Jun17).

All hands meeting

Through the STFC MICE-UK Oversight Committee and communications from STFC to the MICE-UK PI (Paul Soler) the STFC has indicated that its financial position is such that, unless there is a change of circumstance, it is unable to fund the upgrade to Step IV by which we proposed to deliver the demonstration of ionization cooling.

In the light of this development, our ongoing discussions with IHEP Protvino on the possibility of executing a cooling demonstration in the U70 accelerator and the change to the run plan for the present User Cycle, I would like to call an “All Hands” meeting at 15:00 (GMT) on Thursday 25May17. We will use the familiar “video conference” arrangements for the meeting; Melissa Uchida will communicate the details shortly. By using the usual VC slot, I hope that many of you will be able to attend.

With best wishes …

Ken

Update 03 April 2017

Added by Kyberd, Paul 2 months ago

Dear All,

Preparation for the next Cycle continues in the MICE Hall. The LH2 installation programme has been proceeding to plan. Over the past week or so there has been much discussion in the LH2 and engineering teams about the possibility of a pressure excursion in the LH2 safety-vacuum space in the (unlikely) event of a failure of one of the LH2 windows. A plan to resolve the issue was agreed last Friday and work is now progressing once more.

A critical element of the LH2 programme is the test of the system with neon. The test will be to condense neon, establish a depth of liquid neon in the absorber and keep the level stable using the LH2 control system. The work now proceeds to deliver the neon test as soon as possible. We will hear more about the status of the preparations at this week’s video conference.

In parallel, preparations have been made for the servicing of the cold-heads in the tracker cryostats. This is now scheduled to take place on Wednesday this week. Other work planned for the shutdown is being executed systematically.

Comments on the “demo paper” have been received from the referees appointed by PRAB. These comments have been addressed by the paper’s authors. I hope that this paper will now be published very soon.

IPAC’17 (14—19 May 2017) will take place in just over 5 weeks in Copenhagen. We have the ambition to take new data to this conference. Chris Rogers has organised an analysis meeting on the 20th April 2017. This will be the opportunity to identify those analysis that are mature enough to be presented at IPAC.

Finally, registration for CM48 in Belgrade (27–29 June 2017) is now open. Please register at the meeting page:

https://indico.cern.ch/event/626707/

With best wishes …

Ken

Update 10 March 2017

Added by Kyberd, Paul 3 months ago

This week has been eventful for MICE. On Tuesday, operations in flip mode with the lithium-hydride absorber ended, the work to install the liquid-hydrogen absorber started and the review of the experiment by the Resource Loaded Schedule Review panel and the MICE Project Board took place. The following is a brief summary of the outcome of the review, an update on the flip-mode data taking and the preparations for the LH2-system install and the announcement of the date of CM48 in Belgrade.

RLSR/MPB review:

The RLSR panel and the MPB met at RAL on the 7th and 8th of March 2017. The papers prepared for the review as well as the presentations may be found on MICEmine at:

http://micewww.pp.rl.ac.uk/projects/governance/wiki/2017-03-07-Review

The talks were well prepared, well delivered and well received. I would like to thank all the speakers and all those who contributed to the preparation of the documentation that, together, convey so well the excellent progress that has been made by the hard work of the collaboration.

In their feedback at the close-out, the reviewers congratulated the collaboration on delivering “… useful and important scientific data”, noted the outstanding commitment of the MICE personnel and welcomed the fact that the collaboration continues to attract new institutes.

There were only a few actions and recommendations. The reviewers recommended “…very strong[ly] and unanimous[ly] …” that the collaboration be allowed to take data in the September/October ISIS User Cycle (2017/02). In addition, the feedback included the recommendation that we maximise the scientific output of the experiment by writing “… as many peer-reviewed publications as possible”. The reviewers recommended that we extend our series of technical publications to include contributions on the superconducting magnets, the RF system and the overall MICE experiment.

The terms of reference for the review did not include the mandate for the reviewers to comment on our proposal to upgrade Step IV to deliver a cooling demonstration with acceleration. However, the reviewers commented that “… the scientific benefits [of such an upgrade] would be significant …”. The first step in bringing the proposed upgrade forward is the submission of a “Statement of Intent” (SOI) to STFC by MICE-UK. Paul Soler has begun to prepare the SOI based on the material presented at CM47 and presented at the review.

Flip-mode operation with LiH:

Data was taken in flip-mode with LiH during the first three weeks of the present ISIS User Cycle (2016/05). The FC had been run up to 165A on its own on the beam-line. It was therefore a disappointment that it quenched at 160A when run up in the presence of the spectrometer solenoids. Despite this set back, the principal factor that limited our ability to take data was ISIS, which suffered from a number of failures that resulted in the loss of approximately one of the three weeks of data taking.

When data-taking ended on Tuesday morning this week, we had taken two of the three lattice settings that had been planned for the flip-mode data-taking period. Paolo Franchini and Melissa Uchida, who were the MOMs for this Cycle, did an excellent job of ensuring that the MICE up-time was excellent when ISIS was in operation. Initial studies indicate that the data quality is good.

Liquid-hydrogen absorber installation and commissioning programme:
The MICE Hall and the liquid-hydrogen teams moved efficiently to begin preparations for the installation of the liquid-hydrogen absorber as soon as data taking ended. The status is that the PRY has been partially disassembled, the FC has been withdrawn, the LiH absorber placed into safe storage and the spectrometer solenoids have been moved to their nominal positions. This will aid the reinstallation of the FC.

The schedule for the liquid-hydrogen system install is tight. Vicky Bayliss and the LH2-team have carefully planned the installation and commissioning programme—including the necessary safety sign-off. So far the programme remains on track,

CM48: Belgrade 27–29 June 2017

At CM47 we decided that the next collaboration meeting will be held in Belgrade. It is a pleasure to confirm that the meeting will take place between the 27th and 29th June 2017. Further details will be circulated in due course.

With best wishes …

Ken

Update 02 December 2016

Added by Kyberd, Paul 3 months ago

Dear All,

The Step IV data-taking programme with the lithium-hydride absorber continues to proceed according to plan. The present User Cycle, 2016/04, is dedicated to the collection of the data that will allow us to study the evolution of normalised transverse emittance in the Step IV lattice. Chris Rogers, Ao Liu and Jaroslaw Pasternak have carefully selected settings that should allow us to observe normalised-emittance reduction, cooling, over a range of input emittance and beam momentum.

To date two of our target of three settings have been recorded. Chis presented the attached plots at yesterday’s video conference. The plots show the distribution of transverse amplitude for samples of the data taken at 140 MeV/c with nominal input emittance of 6 mm and 10 mm. Good muons are defined to be those that pass through the whole apparatus, i.e. these muons are measured in both the upstream and in the downstream trackers. The green and blue histograms show the distribution of transverse amplitude measured in the upstream and the downstream tracker respectively. The red histograms show the distribution of transverse amplitude measured in the upstream tracker for those muons that are not observed after they have left the upstream tracking volume.

The red histogram indicates that transmission losses are dominated by muons with a large transverse amplitude upstream of the absorber. If the combination of energy loss in the LiH and the optics of the channel are such as to produce ionization cooling, then the population at low transverse amplitude measured in the downstream tracker should be larger than that measured in the upstream tracker. Looking closely at the plots, you can see an indication of such an effect in the (6,140) data. In the (10,140) data, an excess of events at low transverse amplitude is observed in the downstream distribution for transverse amplitudes below ~20 mm. As Chris says on his slide; MICE cools!

This is really an excellent start! We continue to take data in the 2016/04 1.2 configuration. On Monday we plan to reconfigure the magnetic channel to take data with a slightly smaller beta at the absorber. Analysis of the data will then allow us to study the cooling effect with lithium hydride.

If data taking continues with the efficiency that we have established over the past two Cycles, we shall be able to enjoy the Christmas break with lots of “data in the bank”!

With very best wishes …

Ken

Update 26 October 2016

Added by Kyberd, Paul 3 months ago

Dear All,

Data taking for the field-on LiH scattering programme started on Monday 24Oct16.

The attached plots show the difference between the momentum reconstructed in the upstream and downstream trackers for 140MeV/c and 170MeV/c beams. The distributions show that the momentum is decreased by roughly 15MeV/c as the beam passes through the absorber. The size of this momentum loss is in line with expectations.

Both spectrometers were ramped to the currents required by the “6.1.4” scattering configuration late on Sunday 25Oct16.

The FC was ramped to current over night Sunday/Monday. During the day on Monday 24Oct16, ISIS recovered from the issues they had experienced in the vacuum and RF systems. MICE began taking data as soon as ISIS was ready, data taking continues.

By tomorrow morning, we anticipate that substantial scattering data sets will have been accumulated at 140MeV/c, 170MeV/c, 200MeV/c and 240MeV/c. Therefore, in preparation for the November/December Cycle, data will be taken tomorrow with various diffuser settings. This data will allow us to verify our understanding of the effect of the diffuser on the beam; an exercise that we have not performed before.

This is a big step for our physics programme and is the result of a lot of hard work by many, particularly over the past two weeks during which the absorber was installed and gremlins in the solenoid power systems were flushed out.

To recognise what has been achieved, and to initiate the discussion of the programme to be carried out in the next User Cycle, you are cordially invited to a MICE end-of-cycle briefing which will be held in R9 at 12:00 on Friday 28Oct16.

With very best wishes …

Ken

Update 24 September 2016

Added by Kyberd, Paul 3 months ago

Please join me in congratulating all involved in bringing the MICE magnets to current in a configuration typical of the settings required to deliver the field-on scattering programme at Step IV. The magnets have been at current since the evening of Thursday 22Sep16 and a full 48-hour soak test is now in progress. On Friday 23Sep16, the the first data was taken in this configuration with 140 MeV and 170 beams, i.e. we’ve started the empty-channel data taking for the field-on scattering programme.

Update 22 August 2016

Added by Kyberd, Paul 3 months ago

Success at the summer conferences:
MICE presentations at the summer conferences, ICHEP and NuFact, were based around the papers that are in preparation. The bases of the presentations were:
• The first reconstruction of emittance using a beam sampled one particle at a time. These data validate the reconstruction technique we plan to use to study cooling;
• The first measurement of multiple Coulomb scattering (MCS) of muons on lithium-hydride (LiH) using field-off data collected earlier this year; and
• The final results of the optimisation of the cooling-demonstration lattice.

The quality of the presentations was uniformly high. Steve Boyd made the parallel-session presentation at ICHEP, while posters were presented by Victoria Blackmore and Tanaz Mohayai. At NuFact the three parallel-session presenters were Francois Drielsma, Ryan Bayes and Yordan Karadzhov. Each of the presenters gave a very professional impression, reflecting well on the months of hard-work that led to the excellent new results.

Successful combined test of SSU and focus coil:
On the 02Sep16 the magnet-string commissioning passed a most significant milestone. The combination SSU was run in combination with the FC at currents representative of those that are required to deliver the solenoid-mode scattering programme. The field in the tracking volume was 3T, the operating point chosen for the Step IV programme.

SSU was run up so that the ECE combination carried 206A, M2 190A and M1 100A. Then, the FC was run to 75A in small current steps. The attached plot shows the ramp of the FC with SSU energised. The blue trace is the FC current, the black trace is the force measured by the load cells on the FC and the magenta trace is the inter-module force calculated using the currents in the various windings and the coil dimensions. This success will allow the magnet and operations teams to plan the preparation for the User Cycle with confidence.

Magnet-string commissioning will continue next week and on to the start of the next ISIS User Cycle on 13Sep16.

Data in the bank:
A substantial quantity of high-quality data with the superconducting-magnet channel in various configurations was taken in the most recent ISIS User Cycle. This data can be used to study the alignment of the magnetic axis of each magnet individually and to make initial studies of the solenoid-mode optics. It is important that these studies are documented and published the results for future reference.

Commissioning and Hall status update:
Progress has been made in identifying issues in the modules that control the “trim current” to the end coils in the solenoids. A test remains to be performed to demonstrate the stability of the end coils in operation. The source of transients observed in the voltage-tap signals of SSU and SSD remains to be understood fully. In the last User Cycle, stable operation of SSU and SSD was achieved with a voltage-tap threshold of 200mV. A control-software update for the spectrometer solenoids has been prepared at Fermilab and will be deployed next week. Further tests to elucidate the source of the transient signals will then be performed.

There has been steady progress on the preparation of the liquid-hydrogen vessel in R9. The absorber bodies have been passed "good to go" and the MLI has been fitted. Pipe-work modifications are complete. Final level-sensor checks have been carried out and tape to reduce the emissivity of the warm surfaces of the magnet have been installed. The end-caps have been mounted on cradles to allow final preparations to be made. The end caps will be fitted towards the end of next week. Evacuation of the module and the "interspace" can then begin.

The upgrade to the Hall water system is now complete; the demineralised and “inhibited” circuits are now separate; the final modifications having been made on Thursday 01sep16.

The Kaiser compressor serving the Linde refrigerator that cools the decay solenoid was serviced. Following the service, air contamination of the helium was observed and a number of small leaks were found. Kaiser engineers made a visit to the Hall on 31Aug16 and identified two leaks. Work continues to establish that the system is leak tight and, if it is not, to address the remaining issues.

The remedial work on the air-conditioning units is complete. The Hall now has five functioning air conditioners. This, together with the decreasing air temperatures as autumn comes to Oxfordshire, is estimated to be sufficient to stabilise the ambient temperature in the Hall.

ISIS User Cycle 2016/03 (13Sep16 — 29Oct16):
Planning for the next User Cycle is now well underway. The run plan includes data-taking to allow detailed studies of the optics of the Step IV solenoid-mode lattice. The priority for the Cycle is empty-absorber data-taking for the Step IV, field-on scattering programme.

Shift sign-up in CHEESE is live and all are encouraged to sign up.

(1-10/112)

Also available in: Atom